A cell has a terminal voltage of 1.35 V when it is connected to a 0.9 ohm external resistor. 

When the resistance is increased to 2.4 ohm, the terminal voltage is 1.44 V. Calculate the 

internal resistance of the cell.


1
Expert's answer
2022-02-07T15:17:23-0500

"E=V+Ir"


"V=IR"

"1.35=I(0.9)"

"I=1.5"


"\implies\>E=1.35+1.5r...." (i)

"1.44=I(2.4)"

"I=0.6"


"E=1.44+0.6r" ........(ii)



Equation (ii) "-" (i)


"0.09-0.9r=0"

"r=0.1\Omega"